OpenStudy (anonymous):

how do you write an equation in point-slope form for the line through the given point with the given slope. (10,-9) m=-2

OpenStudy (anonymous):

We know \((x_0,y_0)=(10,-9),m=-1\). Just put it into the point slope formula.\[ y-y_0=m(x-x_0) \]

OpenStudy (anonymous):

Another thing you can do is: \[ y_0=mx_0+b \]Solve for \(b\), and put it (along with \(m\)) back into \[ y=mx+b \]

OpenStudy (anonymous):

so.. y-(-9) =-1(x-10)?

OpenStudy (anonymous):

\(m=-2, m\neq -1\) You need to simplify as well.